OpenStudy (anonymous):

Which set of side lengths could form a triangle similar to a triangle with side lengths 11, 15, and 22? A 22, 26, 33 B 22, 30, 42 C 33, 44, 66 D 33, 45, 66

OpenStudy (anonymous):

All 3 sides have to be in proportion, that is, the same exact number is multiplied to each side to get the corresponding side in the new triangle. Just look at the before and after of each side and see if the same number is being used as the multiplier to get the new number.

OpenStudy (anonymous):

is it C?

OpenStudy (anonymous):

So, go through each option one by one. If you start with the third option, we see that 11 has to be multiplied by 3 to get to 33. Is 3 x 15 = 44?

OpenStudy (anonymous):

It is "d" because 3 x 11 = 33 3 x 15 = 45 3 x 22 = 66

OpenStudy (anonymous):

All good now, @stepennd7 ?

OpenStudy (anonymous):

Yes, THANKS! @tcarroll010

OpenStudy (anonymous):

It id option D becausse 33/11=3, 45/15=3 and 66/22= 3 Here the ratios of all the sides are same therefore option D is correct.

OpenStudy (anonymous):

Good luck to you in all of your studies, @stepennd7